Carbon Polarimeter Status

26 Feb 2008: We lost two more targets over the weekend.  One vertical Blue and one vertical in Yellow. Here is a spreadsheet of the target composition and readiness.

01 Nov 2007: Blue has a complete set of targets, and all new silicon detectors.  New target drive mechanism is installed.  Vacuum has be reestablished in Blue, and a calibration run taken.  All channels appear to be functioning well, except for one channel that is probably OK/ but does not see the alpha source, and a second channel that appears to be dead.  Yellow has a full complement of targets, but the old detectors were left in.  Vacuum has be reestablished, and all channels test out OK.  At this point the HV bias has been locked out and the preamp boxes have been removed to improve cooling.  After the boxes have ventilation holes made, they will be replaced.  No schedule for this yet.

27 Sept 2007:  The RHIC vacuum chamber was let up to air in order to install a new target drive mechanism for the rotary positioning system.  The new drive should have less back-lash (hysteresis) and function more reliably.  Work on the drive system will take ~10 days, with a plan to install all new detectors (11 + the detector that was installed but never used last year) and targets, even though the existing targets may be in good shape to operate.. 


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528
Home Phone: 631-878-9020